OpenStudy (anonymous):

how would you change 3x-5y=10 to y=mx=b

OpenStudy (anonymous):

First bring 3x over to the right side of the equation. -5y=-3x+10 (positive becomes negative when you bring it over) Then, bring the negative sign over to the right. 5y=3x-10 Finally, bring 5 over to the right side, where it becomes a division (since 5y=5*y) so, y=3x/5-10/5 and simplify... y=3x/5-2